Charges q1, q2 and q3 are placed at A, C and P, are isosceles triangle vertices and B is the middle point of A and C , and q1 =q2=-q3=2μC. determine the magnitude and the direction of the election field at point P. (P is the top vertices)
1
Expert's answer
2020-03-04T10:07:36-0500

Draw the situation:



Determine the magnitude of the electric field at point P, which is the top vertex. First, find the field created by charge 3:


"E=\frac{kq_3}{r^2}=\frac{kq_3}{0^2}=\infty."

Therefore, no matter what other field components are, the magnitude of the field at P is infinity because the distance between the charge and the point P is zero.

The direction of the field at point P cannot be defined because the field lines are directed from the surrounding space to the third charge.